OpenStudy (anonymous):

how many molecules are in 72.45g CCl4 ?

OpenStudy (anonymous):

\[72.45gCCl4*\frac{ 1 mol cCl4}{ 154gCCl4? }*\frac{ 6.022*10^{23}molecules }{ 1molCCl4 }\] answer= 2.833*10^23 molecules